# Generalist Forecasting with Frozen Video Models via Latent Diffusion

### v1

@@ 4.1 Downstream Tasks and Datasets
We center our evaluation on downstream forecasting tasks designed to span multiple levels of semantic abstraction—from raw pixels to high-level object bounding boxes. This diversity allows us to probe how well frozen video representations support different kinds of future prediction and identify where video models generalize—and where they fail. We visualize each of these tasks in Figure 2(c).
Pixels. We evaluate models on the task of forecasting future RGB frames in ScanNet Dai et al. (2017). While forecasting in pixel space is highly challenging and high-dimensional, it tests low-level generative fidelity and temporal coherence. Pixel forecasting captures fine-grained dynamics but is often sensitive to misalignment or visual ambiguity. We measure pixel accuracy using PSNR.
@@
Point Tracks. Forecasting the trajectories of dense visual features or tracked keypoints in the Perception Test dataset Pătrăucean et al. (2023). Point tracks offer a structured yet fine-grained measure of temporal consistency and motion understanding. Because the same points persist over time, they provide a strong signal for evaluating both representation quality and future modeling. We report Average Jaccard (Doersch et al. (2022)).
Object Bounding Boxes. Forecasting future object locations as bounding boxes focuses on semantic-level understanding of object motion and interaction. This task tests whether representations capture object permanence, affordance, and dynamics—crucial for robotics or autonomous driving applications. We report Mean Intersection over Union (IoU) on the Open Waymo dataset Sun et al. (2020).
@@ 4.2 Benchmarked Models
We benchmark a set of the highest performing and largest image and video models available. See Table 1 for the model and pretrainig specs for all models under consideration.
Image models. We benchmark SigLIP-2B Zhai et al. (2023), a 2B-parameter vision transformer trained on image–text pairs using a contrastive binary classification objective, and DINOv2 Oquab et al. (2023), a 303M-parameter vision transformer trained purely on images using a self-distillation loss without any language supervision. Since these models are not natively trained on video, we follow Carreira et al. (2024) and append learnable temporal positional embeddings to their output features. This modification enables fair comparison with video models by allowing the readout heads to exploit temporal structure when trained on top of the frozen embeddings.
Video models. We evaluate two categories of video models. The first group consists of models trained using masking-based self-supervised objectives. VideoMAE Tong et al. (2022), VideoMAEv2 Wang et al. (2023), and 4DS-e Carreira et al. (2024) are trained to reconstruct masked pixels, while V-JEPA Bardes et al. (2024) uses a feature reconstruction loss based on predictions from a teacher network. VideoPrism Zhao et al. (2024) incorporates language supervision through contrastive learning between video and text during pretraining, followed by a second stage that applies a masked reconstruction loss on video. The second group includes W.A.L.T. Gupta et al. (2024), a video synthesis model trained jointly for frame prediction by conditioning on a past-frames-based signal with a probability $p_{\text{fp}}$. Since W.A.L.T is equipped with native forecasting capabilities, we compare with a pipeline, referred to as N-W.A.L.T, where features are extracted from intermediate layers of the frozen W.A.L.T model by running a single forward pass using the past-frames-based conditioning signal. For N-W.A.L.T we do not need a diffusion model and the predicted features are directly passed to the lightweight readout heads to decode task-specific outputs. We use the same layers for feature extraction as in Vélez et al. (2025).
Model Type Variant Resolution Params Training Time
DINOv2 Oquab et al. (2023) I DINOV2_L/14 224 303M 637.5k steps
SigLIP Zhai et al. (2023) I/L PaLI3-SigLIP-G-opt/14 224 2B 107k steps
VideoPrism Zhao et al. (2024) V/M/L v_giant 288 1.1B 500k steps
VJEPA Bardes et al. (2024) V/M L 224 300M 90k steps
VideoMAE Tong et al. (2022) V/M H 224 600M 1600 epochs
VideoMAEv2 Wang et al. (2023) V/M g 224 1B 1200 epochs
4DS-h Carreira et al. (2024) V/M h 224 639M 488,282 steps
4DS-e Carreira et al. (2024) V/M e 224 4B 488,282 steps
W.A.L.T Gupta et al. (2024) V/S/L 419M 128 419M 528k steps
@@ 4.3 Implementation Details
Readout head and diffusion training. The training methodology for the task-specific readout heads is the same as in Carreira et al. (2024). Readout heads are attention based and trained with L2 error for pixels and depth. For point tracks, a weighted sum of Huber loss of positions and cross entropy over visibility and uncertainty is used. For box tracking, L2 loss between the labeled box coordinates and predicted position is used. Because the base video model latents are frozen, the forecasting diffusion model and the readout head can be trained simultaneously. We found that layer normalization Ba et al. (2016) of the frozen latents is extremely important for forecasting performance. We train for 40k iterations at a batch size of 32 or an equivalent 160k iterations for a batch size of 4 for the memory intensive SigLIP. Aggregating across all experiments, we utilize approximately 144 days worth of tpu-v5 and v6 chips.
Evaluation protocol. For pixels and depth, we use the standard train and validation split on ScanNet Dai et al. (2017). For point tracks, we train in the Kubric movie dataset Greff et al. (2022) and test on the Perception Test dataset Pătrăucean et al. (2023). On box tracking, we use the train and validation split of the Waymo Open Box dataset Sun et al. (2020). For all forecasting tasks we take as context 4 frames and forecast the next 12 frames in all experiments. We sample the diffusion model 10 times per example during evaluation.
W.A.L.T setup. We utilized W.A.L.T., a text-to-video diffusion model, as a frozen encoder, probing its intermediate layers in a setup similar to Vélez et al. (2025). W.A.L.T. is designed to process 17 video frames, tokenizing them into five latent representations: one for the initial frame and four for the subsequent 16 frames. To maintain consistency with other models in this study, we sampled 16 frames, duplicated the first frame, and simulated the forward diffusion process by adding noise at timestep $t$. Instead of the complete multi-step generative process, the visual representation is obtained by a single forward pass through the denoiser, utilizing a null text embedding. During the single pass, the intermediate representations are extracted, discarding the initial latent representation. We use the same layers for feature extraction as in Vélez et al. (2025).
@@ 5 Results
Figure 3: Forecasting per-example metric results. We evaluate forecasting on pixels (PSNR), point tracks (Jaccard Distance), bounding box tracks (IoU), and depth maps (Mean Absolute Relative Error) using 10 samples per example. For reference, we also report perception performance on each task. Given the stochastic nature of forecasting, we report the mean and maximum/minimum performance across samples. This reveals differences in sample quality not captured by the mean alone—some models exhibit similar averages but differ significantly in their best-case outputs, highlighting variation in their predictive distributions. Overall, we observe that stronger perception models tend to yield better forecasting performance.
Task Pixels Depth Points Boxes
Dataset ScanNet ScanNet Perc. Test Waymo
@@
DINOv2 82.3 2.2 626.75 7.3 1.9 0.038 3.08 0.044
SigLIP 182.74 0.5 872.09 3.0 3.0 0.038 3.22 0.05
VideoPrism 83.44 4.2 948.14 5.3 0.8 0.039 2.72 0.045
VJEPA 61.46 2.2 710.5 2.9 0.63 0.039 2.85 0.048
VideoMAE 33.76 5.8 643.82 5.4 0.55 0.039 2.62 0.045
VideoMAEv2 45.57 3.7 652.53 4.0 0.74 0.039 2.92 0.048
4DS-h 32.52 7.0 616.23 5.0 7.88 0.039 3.24 0.054
4DS-e 32.26 6.4 599.43 5.6 0.68 0.039 1.87 0.036
W.A.L.T. 500M 8.60 4.96 222.92 3.2 1.34 0.039 2.47 0.040
N-WALT 500M 10.9 3.5 263.96 1.7 1.3 0.038 2.31 0.038
@@
Table 2: Distributional alignment of forecasted futures. We report Fréchet Distance (lower is better) and the variance of fitted Gaussian distributions for each metric, comparing the ground truth distribution to the model’s sampled forecasts. Ideally, the forecasted variance should closely match the ground truth. Results show that stronger perception models produce forecasts with distributions more aligned to the data, reinforcing trends observed in the per-example metrics (Fig. 3).
Quantitative. We present our experimental results in Figure 3, reporting per-example forecasting metrics alongside perception performance, and Table 2, summarizing distribution metrics. Overall, we observe a strong correlation between per-example metrics and the distributional alignment of predicted futures with ground truth, as captured by FD and variance. However, notable discrepancies emerge between the two evaluation paradigms. For example, in depth forecasting, per-example metrics suggest that models like VideoPrism perform comparably to stronger models, while FD reveals significant underperformance—highlighting how small-sample evaluations can obscure poor distributional alignment.
Our central finding is that forecasting performance tends to correlate with perception performance, particularly for masked modeling approaches. Among these, 4DS-e consistently outperforms all other models on FD across most tasks, in line with its strong perception results—except in point track forecasting, where VideoMAE performs best. Interestingly, VideoMAEv2 underperforms its predecessor across nearly all metrics, while VideoMAE (v1) shows strong results in bounding box forecasting (per-example) and point tracks (FD), despite its smaller model size.
Models pretrained exclusively on image-based objectives, such as DINOv2 and SigLIP, perform poorly across most tasks, reinforcing the importance of temporal supervision during pretraining. Similarly, language-augmented models like SigLIP and VideoPrism, which were trained only on perception-style tasks, also lag behind.
@@
When evaluating masked representation-learning models of similar size (VideoMAE and 4DS-H) against W.A.L.T., FD reveals that W.A.L.T. significantly excels at pixel and depth forecasting—tasks closely aligned with its training objective. W.A.L.T. also outperforms 4DS-H in both point and bounding box forecasting, while performing comparably to VideoMAE in these tasks. This outcome does not align with W.A.L.T.’s lower perception performance when benchmarked against these two models for depth prediction and object tracking. It is worth noting that N-WALT does not exhibit the same performance. It excels at pixel forecasting since it was trained with a frame prediction objective, but it performs poorly on the other tasks, which indicates that the learned features are not useful across different levels of abstraction.
@@
We find that all models on most tasks, with the exception of point tracking, struggle to approach the variance of the ground truth datasets. Even though W.A.L.T performs relatively well on pixel forecasting with respect to the Frechet Distance metric, it still does not capture the full extent of the underlying variance in pixel space. The variance disparity is especially apparent with depth forecasting; this suggests that current models particularly struggle to model visual information relevant to this particular domain.
@@ 6 Discussion
Our central finding is that forecasting performance in frozen pretrained video models closely tracks their perception performance. Models that excel at perception tasks—especially those trained with temporally coherent video data—also tend to perform well when repurposed for forecasting. In contrast, language supervision alone does not appear to improve forecasting ability, underscoring the importance of temporal visual learning for anticipating future states.
As expected, W.A.L.T., a video synthesis model explicitly trained to generate future frames, significantly outperforms masked video models on low-level forecasting tasks such as pixel and depth prediction. However, W.A.L.T.’s forecasting strength is mixed for mid-level structured tasks like point tracks and object bounding boxes when comparing masked models of similar size. This varied capability is surprising given W.A.L.T.’s perception performance where it generally lags behind with the exception of pixel prediction. One possible explanation lies in W.A.L.T.’s training regimen: its objective is dominated by text-to-video generation (90% of the time), with direct forecasting used in only 10% of training. This imbalance may limit its ability to internalize the structured, temporally consistent representations required for higher-level forecasting tasks.
Limitations. Our forecasting framework has a few key limitations. First, the datasets studied—while diverse—often lack complex or ambiguous motion, limiting the generality of the tasks. Second, while diffusion provides a unified and expressive forecasting mechanism, it is computationally expensive and sensitive to sample count. Third, forecasting on frozen perception features may impose representational mismatches, especially for tasks requiring fine-grained temporal dynamics. Lastly, we use video models with 16-frame contexts, which restricts our ability to assess long-horizon forecasting. In the short-horizon world, forecasting performance strongly correlates with perception, but this may not be true for longer horizons, an investigation we leave for future.
Conclusion. We introduced a unified forecasting framework utilizing frozen video models across diverse tasks and levels of abstraction. Our results reveal that forecasting performance strongly correlates with perception performance, especially in models trained with video supervision. Generative models like W.A.L.T. meet or exceed the performance of mask-based models as is expected from their generative training objective. These findings position forecasting not just as a test of generative fidelity, but as a powerful lens on representation quality—paving the way for future research into temporally generalizable video understanding.

### v2

@@ 4.1 Downstream Tasks and Datasets
We center our evaluation on downstream forecasting tasks designed to span multiple levels of semantic abstraction, from raw pixels to high-level object bounding boxes. This diversity allows us to probe how well frozen video representations support different kinds of future prediction and identify where video models generalize, and where they fail. We visualize each of these tasks in Figure 1(c).
Pixels. We evaluate models on the task of forecasting future RGB frames in ScanNet [dai2017scannet]. While forecasting in pixel space is highly challenging and high-dimensional, it tests low-level generative fidelity and temporal coherence. Pixel forecasting captures fine-grained dynamics but is often sensitive to misalignment or visual ambiguity. We measure pixel accuracy using PSNR.
@@
Point Tracks. Forecasting the trajectories of dense visual features or tracked keypoints in the Perception Test dataset [patraucean2023perception]. Point tracks offer a structured yet fine-grained measure of temporal consistency and motion understanding. Because the same points persist over time, they provide a strong signal for evaluating both representation quality and future modeling. We report Average Jaccard ([doersch2022tap]).
Object Bounding Boxes. Forecasting future object locations as bounding boxes focuses on semantic-level understanding of object motion and interaction. This task tests whether representations capture object permanence, affordance, and dynamics—crucial for robotics or autonomous driving applications. We report Mean Intersection over Union (IoU) on the Open Waymo dataset [Sun_2020_CVPR].
@@ 4.2 Benchmarked Models
We benchmark a set of the highest performing and largest image and video models available. See the supplementary material for the model and pretrainig specs for all models under consideration.
Image models. We benchmark SigLIP-2B [zhai2023siglip], a 2B-parameter vision transformer trained on image–text pairs using a contrastive binary classification objective, and DINOv2 [oquab2023dinov2], a 303M-parameter vision transformer trained purely on images using a self-distillation loss without any language supervision. Since these models are not natively trained on video, we follow [carreira2024scaling] and append learnable temporal positional embeddings to their output features. This modification enables fair comparison with video models by allowing the readout heads to exploit temporal structure when trained on top of the frozen embeddings.
Video models. We evaluate two categories of video models. The first group consists of models trained using masking-based self-supervised objectives. VideoMAE [tong2022videomae], VideoMAEv2 [wang2023videomaev2], and 4DS-e [carreira2024scaling] are trained to reconstruct masked pixels, while V-JEPA [bardes2024vjepa] uses a feature reconstruction loss based on predictions from a teacher network. VideoPrism [zhao2024videoprism] incorporates language supervision through contrastive learning between video and text during pretraining, followed by a second stage that applies a masked reconstruction loss on video. The second group includes WALT [gupta2024walt], a video synthesis model trained jointly for frame prediction by conditioning on a past-frames-based signal with a probability $p_{\text{fp}}=0.1$. We leverage this built-in capability in a pipeline referred to as Native WALT (N-WALT). N-WALT is not a new model; it is simply the pretrained WALT model used exclusively in its forecasting mode. For this pipeline, a single forward pass is performed with the past-frames-based conditioning signal to extract predictive features from the model intermediate layers. These features are then directly decoded by lightweight readout heads to produce task-specific outputs, thereby obviating the need for a separate diffusion model. We use the same layers for feature extraction as in [velez2025walt4ds].
@@ 4.3 Implementation Details
Forecasting diffusion model and readout head. Our diffusion implementation uses DDIM [song2021denoising] and incorporates a cosine schedule [nichol2021improved]. The underlying backbone denoiser is a vanilla 5-layer transformer. Each transformer layer employs multi-headed attention with 8 heads, utilizing 1024 total dimensions for queries, keys, and values, alongside a 2048-dimension hidden layer for the MLP. The training objective for the diffusion model minimizes the mean squared error between the denoised output of the model and the original latents, computed from a given video encoder. The diffusion model architecture is consistent across all of the underlying video models.
The training methodology for the task-specific readout heads is the same as in [carreira2024scaling]. Readout heads are attention based and trained with L2 error for pixels and depth. For point tracks, a weighted sum of Huber loss of positions and cross entropy over visibility and uncertainty is used. For box tracking, L2 loss between the labeled box coordinates and predicted position is used.
Because the base video model latents are frozen, the forecasting diffusion model and the readout head can be trained simultaneously. We found that layer normalization [ba2016layernormalization] of the frozen latents is extremely important for forecasting performance. We train for 40k iterations at a batch size of 32 or an equivalent 160k iterations for a batch size of 4 for the memory intensive SigLIP. Aggregating across all experiments, we utilize approximately 144 days worth of tpu-v5 and v6 chips.
Evaluation protocol. For pixels and depth, we use the standard train and validation split on ScanNet [dai2017scannet]. For point tracks, we train in the Kubric movie dataset [greff2021kubric] and test on the Perception Test dataset [patraucean2023perception]. On box tracking, we use the train and validation split of the Waymo Open Box dataset [Sun_2020_CVPR]. For all forecasting tasks we take as context 4 frames and forecast the next 12 frames in all experiments. We sample the diffusion model 10 times per example during evaluation. We use TPUs ranging from 4-32 chips (v5p and v6e) and report total training time for the experiments in Table 7 in the appendix.
WALT setup. We utilized WALT, a text-to-video diffusion model, as a frozen encoder, probing its intermediate layers in a setup similar to [velez2025walt4ds]. WALT is designed to process 17 video frames, tokenizing them into five latent representations: one for the initial frame and four for the subsequent 16 frames. To maintain consistency with other models in this study, we sampled 16 frames, duplicated the first frame, and simulated the forward diffusion process by adding noise at timestep $t$. Instead of the complete multi-step generative process, the visual representation is obtained by a single forward pass through the denoiser, utilizing a null text embedding. During the single pass, the intermediate representations are extracted, discarding the initial latent representation. We use the same layers for feature extraction as in [velez2025walt4ds].
@@ 5 Results
The need for a stochastic evaluation. To demonstrate the need for modeling the stochasticity of future events, we perform an ablation of our proposed diffusion forecasting module and compare it to regression-based forecasting in Table 1. We find that while regression optimizes the traditional mean-based metric, this does not account for the inherent stochasticity of forecasting. When considering metrics such as Fréchet Distance or Best-of-N that take this stochasticity into account, we find that diffusion models generally outperform the regression baselines in most cases.
Model Pixels Depth Point Tracks Box Tracks
4DS-e Reg. 18.96 18.96 46.61 0.188 0.188 694.18 0.59 0.59 0.00070 0.58 0.58 2.26
4DS-e 19.89 22.03 30.95 0.1937 0.096 533.0 0.58 0.61 0.00068 0.56 0.66 1.87
WALT Reg. 21.69 21.69 14.4 0.2196 0.2196 209.86 0.61 0.61 0.00140 0.54 0.54 2.44
WALT 500M 20.4 22.55 5.46 0.230 0.138 210.1 0.64 0.68 0.00134 0.50 0.58 2.47
@@
Figure 2: Forecasting per-example metric results. We evaluate forecasting on pixels (PSNR), point tracks (Jaccard Distance), bounding box tracks (IoU), and depth maps (Mean Absolute Relative Error) using 10 samples per example. The colored bars represent the best of n metrics for a particular task. Blue represents tasks where higher performance is better, while gold represents a metric where lower is better. We also report perception performance on each task as gray bars. Given the stochastic nature of forecasting, we report the mean (as whisker plot with standard deviation) and maximum/minimum performance (colored bars) across samples. This reveals differences in sample quality not captured by the mean alone—some models exhibit similar averages but differ significantly in their best-case outputs, highlighting variation in their predictive distributions. Overall, we observe that stronger perception models tend to yield better forecasting performance. However, with the exception of box tracking, the best model in forecasting is never the best in perception.
Task (Dataset) Pixels (ScanNet) Depth (ScanNet) Points (Perc. Test) Boxes (Waymo)
@@
DINOv2 62.97 3.1 588.36 8.9 1.9 0.038 3.08 0.044
SigLIP 203.32 0.5 849.18 2.4 3.0 0.038 3.22 0.05
VideoPrism 70.30 6 882.11 7.3 0.8 0.039 2.72 0.045
VJEPA 37.08 5.1 558.28 7.7 0.63 0.039 2.85 0.048
VideoMAE 28.14 7 547.40 9.2 0.55 0.039 2.62 0.045
VideoMAEv2 33.75 5.9 578.51 7.7 0.74 0.039 2.92 0.048
4DS-h 28.29 10.0 555.11 8.5 7.88 0.039 3.24 0.054
4DS-e 30.95 6.2 533.00 11.0 0.68 0.039 1.87 0.036
WALT 500M 5.46 6.9 210.10 5.6 1.34 0.039 2.47 0.040
N-WALT 500M 6.55 5.4 217.8 4.3 1.39 0.038 3.23 0.055
@@
Table 2: Distributional alignment of forecast futures. We report Fréchet Distance (lower is better) and the variance of fitted Gaussian distributions for each metric, comparing the ground truth distribution to the model’s sampled forecasts. Ideally, the forecast variance should closely match the ground truth. Results show that stronger perception models produce forecasts with distributions more aligned to the data, reinforcing trends observed in the per-example metrics (Fig. 2).
Forecasting mostly correlates with perception. Overall, we observe a strong correlation between per-example metrics and the perception performance in Figure 2 where forecasting performance is displayed alongside perception performance. A more nuanced picture emerges when looking at the table in more detail. Here we find that, with the exception of box tracking, the best model in perception for a given task is not the best model for forecasting. At higher levels of performance, it appears that the dynamics of certain representations are inherently easier to model than others even if other representations contain more information relevant to the downstream task. In addition, we find that none of the representations we tested are universally optimal for both forecasting and perception on all tasks. We note that in Figure 2, Walt is the best on pixel tracks, Native WALT on pixels, 4DS-e on box tracks, and DinoV2 on Depth. Similarly, WALT excels on pixels and depth while VideoMAE excels on pixel tracks, and 4DS-e excels on box tracks. These results imply that specific models may be optimal for forecasting but not for perception on a given task.
@@
Synthesis models like WALT achieve forecasting performance on par with or better than models trained with mask-based objectives. WALT significantly excels at pixel and depth forecasting—tasks closely aligned with its training objective—as revealed by FD when evaluating against masked representation-learning models of similar size (VideoMAE and 4DS-h). WALT also outperforms 4DS-h in both point and bounding box forecasting, while performing comparably to VideoMAE in these tasks. This outcome does not align with the lower perception performance of WALT when benchmarked against these two models for depth prediction and object tracking. It is worth noting that N-WALT does not exhibit the same performance. Since it was trained with a frame prediction objective and is conditioned on past frames, it excels at pixel forecasting, effectively capturing low-level spatiotemporal dynamics. However, it underperforms in other tasks that require higher-level semantic understanding, such as point and box tracking. This performance disparity reveals a fundamental limitation of the pixel prediction objective, suggesting that the learned features are not truly generalizable.
Vision-language contrastive training does not result in better forecasting. In Table 2, we find that models with vision-language contrastive supervision like SigLIP and VideoPrism, which were trained only on perception-style tasks, lag behind.
Video backbones outperform image ones. Notably, we find that models pretrained exclusively on image-based objectives, such as DINOv2 and SigLIP, perform poorly across most tasks, reinforcing the importance of temporal supervision during pretraining. This is contrary to the widespread belief that image-based models are sufficient [baldassarre2025back]. We note that the predictor model in  [baldassarre2025back] also uses past temporal context to forecast the future. However, [baldassarre2025back] claims that the base visual representation on top of which this temporally-aware predictor operates can simply be image-based, while we demonstrate experimentally that base models trained on video outperform those trained on images alone. One notable exceptions to this finding is DINO which excels in depth forecasting, perhaps because ScanNet, the dataset we use in our evaluation for depth has limited motion and thus requires less dynamics information for predicting the future.
Per-example vs. distribution-level metrics. To compare our two proposed metric approaches, we first observe the per-example forecasting metrics in Figure 2. Unsurprisingly N-WALT is the strongest in forecasting pixels, given it was directly trained to do so. However, for depth forecasting, DinoV2 seems to exhibit the best per-example results. WALT with the trained diffusion head, but not N-WALT, is the best on forecasting point tracks. Interestingly, VideoMAEv2 underperforms its predecessor across nearly all metrics, while VideoMAE (v1) shows strong results in bounding box forecasting (per-example) and point tracks (FD), despite its smaller model size.
We next turn to the distribution level metrics in Table 2. Overall, we observe a strong correlation between the per-example metrics in Figure 2 and the distributional alignment of predicted futures with ground truth, as captured by FD and variance. However, we also find notable discrepancies emerge between the two forecasting evaluation paradigms. For instance, in depth forecasting, WALT performs strongly in terms of Frechet Distance, but DinoV2 seems to exhibit better per-example results. Similarly WALT excels on per-example metrics for point tracking, yet VideoMAE is better in terms of Frechet Distance on this task. This discrepancy highlights how small-sample evaluations can obscure poor distributional alignment.
@@
We find that all models on most tasks, with the exception of point tracking, struggle to approach the variance of the ground truth datasets. Even though WALT performs relatively well on pixel forecasting with respect to the Frechet Distance metric, it still does not capture the full extent of the underlying variance in pixel space. The variance disparity is especially apparent with depth forecasting; this suggests that current models particularly struggle to model visual information relevant to this particular domain.
The N-WALT results reported in Table 2 and Figure 2 reflect performance at the specific noise levels that yielded the best metric values. The metrics across various noise levels are presented in Tables 5 and 6 in the appendix, with the results suggesting that maintaining low noise levels is critical for achieving the best results across tasks.
@@
Ablations. We ablated the effect of the readout head capacity as well as the number of samples taken from the diffusion model. These results are shown in the appendix Tables 9 and  8 respectively. We find that our results do not change dramatically with respect to these variables.
@@ 6 Discussion
We proposed a unified evaluation framework of frozen vision backbone models in forecasting tasks. Our central findings are first, that forecasting performance in frozen pretrained video models closely tracks their perception performance up to a point. At higher levels of performance, this association starts to break down. Second, as expected, WALT, a video synthesis model explicitly trained to generate future frames, significantly outperforms masked video models on low-level forecasting tasks such as pixel and depth prediction. However, WALT’s forecasting strength is mixed for mid-level structured tasks like point tracks and object bounding boxes when comparing masked models of similar size. Third, language supervision alone does not appear to improve forecasting ability, underscoring the importance of temporal visual learning for anticipating future states. Lastly, our results clearly show that video backbone models outperform their image-based counterparts in supporting future-forecasting tasks.
Limitations. Our forecasting evaluation framework has a few key limitations. First, the datasets studied, while diverse, often lack complex or ambiguous motion, limiting the generality of the tasks. Second, the introduction of a diffusion model is a potential confounding variable in evaluating the forecasting capabilities of frozen models. To mitigate this risk, we opt to use a simple “vanilla” transformer to provide a standardized and simple forecasting module while preventing an overly powerful, task-specific forecaster from compensating for weaknesses in a backbone’s representations. We note that while diffusion provides a unified and expressive forecasting mechanism, it is computationally expensive and sensitive to sample count. Third, forecasting on frozen perception features may impose representational mismatches, especially for tasks requiring fine-grained temporal dynamics. Fourth, it is difficult to perfectly control for factors such as the number of model parameters, resolution of feature spaces, and the pre-training data volume for each frozen pretrained model. Lastly, we use video models with 16-frame contexts, which restricts our ability to assess long-horizon forecasting. That said, our evaluation framework is generic in that it will apply in exactly the same way to longer-context models, once they become available.